OpenStudy (anonymous):

is x=x+3 an inconsistent equation? why or why not?

ganeshie8 (ganeshie8):

Are there any values for `x` that satisfy the given equation ?

OpenStudy (anonymous):

inconsistent equation (roughly) means that there are no x-solutions that satisdy both of the statements. When you say, x=x+3, that is just like saying y=x+3 y=x you can see though, that no matter what you plug for x into the equations, their y solutions won't ever be same. lets just take a look at the graphs, https://www.desmos.com/calculator/sn2fm04xmt See? they don't intersect, and in order for a system of equations to have a solution yo have to have two of the graphs intersect with weach other.